OpenStudy (anonymous):

differentiate the function g(r)=r^2(ln(2r+1))

OpenStudy (anonymous):

\[g \left( r \right)=r^2\left( \ln \left( 2r+1 \right) \right)\]like this?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

so use the product rule

OpenStudy (anonymous):

could you break it down for me? Having trouble breaking it apart

OpenStudy (anonymous):

I need to know how to solve the problem

OpenStudy (anonymous):

u = r^2 v = ln(2r+1) f(r)=u*v => f'(r) = u'*v + u*v'

OpenStudy (anonymous):

what is u'?

OpenStudy (anonymous):

2r

OpenStudy (anonymous):

and what is v'?

OpenStudy (anonymous):

ill take a guess at ln(2+0)?

OpenStudy (anonymous):

nope... \[\frac{ d }{ dr }\ln \left( f \left( r \right) \right)=\frac{ f'\left( r \right) }{ f \left( r \right) }\]

OpenStudy (anonymous):

ln(2)/ln(2r+1)?

OpenStudy (anonymous):

or just 2/2r+1?

OpenStudy (anonymous):

there you go!

OpenStudy (anonymous):

2r+1

OpenStudy (anonymous):

yeah, the second one is correct...

OpenStudy (anonymous):

\[v'=\frac{ 2 }{ 2r+1 }\]

OpenStudy (anonymous):

so v' is 1/r+1?

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

now put it all together

OpenStudy (anonymous):

2r*ln(2r+1)+r^2*(2)/(2r+1)

OpenStudy (anonymous):

looks good!

OpenStudy (anonymous):

\[g'\left( r \right)=2r \cdot \ln \left( 2r+1 \right)+\frac{ 2r^2 }{ 2r+1 }\]

OpenStudy (anonymous):

excellent! thank you is this the best /simpliest form?

OpenStudy (anonymous):

i guess you could factor...\[g'\left( r \right)=2r \left[ \ln \left( 2r+1 \right) +\frac{ r }{ 2r+1 }\right]\]

OpenStudy (anonymous):

thanks!